Assume the following information for a bank quoting on spot exchange rates: Exchange rate of Singapore dollar in USD = 0.32 USD/SGD Exchange rate of pound in USD = 1.50 USD/GBP Exchange rate of pound in Singapore dollars = 4.50 SGD/GBP Based on the information given, as you and others perform triangular arbitrage, what should logically happen to the spot exchange rates?

Choose all that apply.

USD/SGD goes up

USD/SGD goes down

USD/GBP goes up

USD/GBP goes down

SGD/GBP goes up

SGD/GBP goes down

OMG/LOL goes sideways
Answer #1

Lets say u have 1 USD
Sell USD and buy SGD=1/0.32 SGD
Sell SGD and buy GBP=1/4.5*1/0.32 GBP
Sell GBP and buy USD=1.5*1/4.5*1/0.32=1.04167 USD

So profit of 0.04167 USD per USD

As people will sell USD and buy SGD USD/SGD goes up
As people will sell SGD and buy GBP SGD/GBP goes up
As people will sell GBP and buy USD USD/GBP goes down

USD/SGD goes up
USD/GBP goes down
SGD/GBP goes up